OpenStudy (anonymous):

What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. –5 m/s2 B. +5 m/s2 C. –25 m/s2 D. +25 m/s2

OpenStudy (anonymous):

I need to know as soon as possible! I will give medals.

OpenStudy (surry99):

use v = vinitial +at

OpenStudy (anonymous):

\[\color{green}{\vec{a} = \frac{\Delta \vec{v}}{\Delta t} = \frac{\vec{v}_f - \vec{v}_i}{\Delta t}}\]

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

The equation ghuczek gave is just saying to take the difference in velocity and divide it by the change in time: initial speed: 25 m/s final speed: rest = 0 m/s initial time: 0 s final time: 5 s acceleration = (final speed - initial speed) / (final time - initial time)

OpenStudy (anonymous):

Thank you guys!!!